OpenStudy (anonymous):

An astronaut on the moon throws a ball vertically upwards. The height, s metres, of the ball, after t seconds, is given by the equation s= 40t + 0.5at^2, where a is a constant. If the ball reaches its maximum height when t = 25, find the value of a.

OpenStudy (anonymous):

a=-1.6 differenciate wrt t ds/dt = 40+ at to maximise s, make the ds/dt zero.. so, a=-40/t=-40/25=-1.6

OpenStudy (anonymous):

a will be -1.6

OpenStudy (anonymous):

THANK YOU!! :D

OpenStudy (anonymous):

could someone explain the process of solving that problem?